Find the corresponding y values for x = -4, -3, -2, -1, 0, 1, 2 if y=x^2+2x-3.
Show your work.

Answer by checkley77(12844)

x = -4, -3, -2, -1, 0, 1, 2 if:
y=x^2+2x-3
-----------------
y=-4^2+2*-4-3
y=16-8-3
y=5
-----------------------------------
y=-3^2+2*-3-3
y=9-6-3
y=0
-----------------------------------
y=-2*2+2*-2-3
y=4-4-3
y=-3
I'll leave the rest for you to practice on.